OpenStudy (anonymous):

The sequence is geometric. What is the next term? 12, 24, 48, 96,... Answer 186 192 288

OpenStudy (yamyam70):

12 + 12 = 24 24 +24 = 48 48 +48 = 96 96 + 96 = ?

OpenStudy (anonymous):

192

OpenStudy (anonymous):

oh, O_o thx!

OpenStudy (yamyam70):

:)

OpenStudy (anonymous):

96*2